OpenStudy (jackierobinson12):

State whether the given measurements determine zero, one, or two triangles. B = 86°, b = 21, c = 18

OpenStudy (acespeedfighter):

42

OpenStudy (acespeedfighter):

actually 1-rrkdm jmkmdkm dndjnendql, ht51gt7897geq 594vge+dqw8wef6 4sda2fg2c xf fwrqsd

OpenStudy (mathmate):

@AceSpeedFighter Wow, can't break that code.

OpenStudy (mathstudent55):



OpenStudy (acespeedfighter):

its easy 23432x312312x0-e3i=-23i0-x0i03ii-0ix9ixiu90ux01u3=-1x04=0=x1 ]

OpenStudy (jackierobinson12):

ok then I would use the law of sines?

OpenStudy (jackierobinson12):

Would it be two triangles?

OpenStudy (acespeedfighter):

yes

OpenStudy (acespeedfighter):

no

OpenStudy (acespeedfighter):

65+42 is like 21

OpenStudy (jackierobinson12):

@mathstudent55 @mathmate ??

OpenStudy (acespeedfighter):

so divide it by 10 and 9

OpenStudy (acespeedfighter):

you get 8

OpenStudy (acespeedfighter):

so its siple

OpenStudy (acespeedfighter):

Times it by the times =you said hello

OpenStudy (mathstudent55):

Use the law of sines to find C. Then use the law of sines again to find a

OpenStudy (jackierobinson12):

b/sinB = c/sinC (sine law) 21/sin86° = 18/sinC sinC = sin86° × (18/21) C = 58.8° or C = 180° - 58.8° C = 58.8° or C = 121.2° (rejected for B + C = 207.2° > 180°)

OpenStudy (jackierobinson12):

right?

OpenStudy (mathmate):

Yes, well-done! :)

OpenStudy (jackierobinson12):

so two or one triangles?

OpenStudy (mathmate):

How many solutions do you have for sin C? That would correspond to the number of possible triangles.

OpenStudy (jackierobinson12):

I think 2

OpenStudy (mathstudent55):

Correct. We have an angle of 86 deg. Angle C is 58.8 or 121.2, but 121.2 is too large bec of angle of 86 deg. That means angle C can have only 1 measure, 58.8, and there is only 1 triangle.

OpenStudy (jackierobinson12):

:) Thank you!

OpenStudy (mathstudent55):

yw

OpenStudy (jackierobinson12):

Can you help with another?

OpenStudy (jackierobinson12):

Solve the triangle. A = 19°, C = 102°, c = 6

OpenStudy (mathmate):

Use the sine rule, sin C/c = sin A /a and find a. Since one of the angles is obtuse, the two remaining angles must be acute, and hence only one possible solution.